OpenStudy (anonymous):

Solve, using linear combination. 3x + y = 4 2x + y = 5

OpenStudy (anonymous):

2*(3x+y=4) then 6x+2y=8 -3*(2x+y=5) then - 6x-3y=-15 combine them -1y=-7 then y=7 2x+y=5 then 2x=5-7 x=-1

OpenStudy (anonymous):

linear combination means adding the lines together. you want to multiply one of the equations so that one of the variables becomes the additive inverse of the other to eliminate it. this method is also called elimination.